# Understanding Oversmoothing in GNNs as Consensus in Opinion Dynamics

### v1

@@ C.2.2 Detail for node classification across different depths
Appendix D Additional Experiments
(a) Graph
(b) $-\ell$
(c) $-\ell$
(d) GraphCON-Tran
(e) BIMP (ours)
Figure 3: Toy simulations Result. Given a fully connected graph shown in (a), we simulate and visualize the dynamics of $-\ell$, $-\ell$, GraphCON-Tran and BIMP using the Euler method. All systems share the same adjacency matrix and initial state. The visualized result indicates all the methods, except BIMP, suffer from the oversmoothing as features for all three nodes converges to the same values.
D.1 The Numerical Simulation Result
In order to illustrate the shortcoming across the exiting methods while our method doesn’t, we simulate a toy dynamics of: $-\ell$, $-\ell$, GraphCON-Tran and BIMP. This visualization helps to understand the theoretical analysis in Section 5. The setting is
$\mathbf{A}=\begin{bmatrix}0&0.43&0.57\\ 0.64&0&0.36\\ 0.70&0.30&0\end{bmatrix},\quad\mathbf{X}(0)=\begin{bmatrix}0.43&0.29&0.61\\ 0.14&0.29&0.37\\ 0.46&0.79&0.20\end{bmatrix}.$ (113)
Particularly, in $-\ell$, all the nodes are ‘trustworthy’ to calculate the source term. In BIMP, we randomly generate the belief matrix $\mathbf{A}_{\mathrm{o}}$ using the same method as for $\mathbf{A}$.
In right side of Figure 3, we present the simulated result for these four methods. We observe that features in $-\ell$ (upper left) quickly converge to the same values. In GraphCON-Tran (lower left), features also converge to the same but at slower speed rate due to its damping term resisting to the changes. Features in $-\ell$ (upper right) remain a fixed difference between nodes after $t=4$, but this difference get ignorable over time, ultimately leading to feature similarity. In contract, BIMP keeps distinct features throughout and mitigates the oversmoothing.
Table 7: Classification accuracy on heterophilic datasets. Classification accuracy on the Texas, Wisconsin and Cornell datasets. Our BIMP model outperforms most of the continuous-depth baselines.
Model Texas Wisconsin Cornell
Homophily level 0.11 0.21 0.30
BIMP (ours) 82.16$\pm$4.06 85.29$\pm$3.42 77.13$\pm$3.38
$-\ell$ 74.59$\pm$5.43 82.75$\pm$3.90 70.00$\pm$6.22
GraphCON-GCN 80.54$\pm$4.49 84.79$\pm$2.51 74.05$\pm$3.24
KuramotoGNN 81.81$\pm$4.36 85.09$\pm$4.42 76.02$\pm$2.77
D.2 Experiment on heterophilic datasets
We have evaluated our model on homophilic datasets in main paper, which holds the assumption that edges tend to connect similar nodes. However, many GNN models struggle with low accuracy on heterophilic datasets, where this assumption no longer holds. We deploy our BIMP model on three heterophilic datasets: Texas, Wisconsin, and Cornell from the CMU WebKB (Craven et al., 1998) project. BIMP demonstrates competitive performance while maintaining low computational complexity.
Table 7 lists the classification accuracy of BIMP and other continuous-depth GNNs on these heterophilic datasets. For all benchmarks, we run 10 fixed splits for each dataset with 20 random seeds for each split on work stations with an Intel Xeon Gold 5220R 24 core CPU, an Nvidia A6000 GPUs, and 256GB of RAM. The hyperparameters for BIMP are searched by Ray Tune process with 200 random trails.
Our BIMP model outperforms baselines in all datasets, illustrating its improved performance on heterophilic datasets.

### v2

@@ E.1 Computational complexity
E.2 Choice of nonlinearity in NOD module
To understand how the choice of nonlinearity in our Nonlinear Opinion Dynamics (NOD) module impacts performance, we experiment with a suite of alternative nonlinearities (softsign, arctan, sigmoid, ReLu and GELU) and linearity (linear). Softsign and arctan satisfy the nonlinearity constraint in the NOD definition (i.e., $S(0)=0,S^{\prime}(0)=1,S^{\prime\prime}(0)\neq 0$), but sigmoid, ReLu and GELU do not. Specifically, sigmoid does not pass through the origin, ReLu is not differentiable, and GELU does not satisfy $S^{\prime}(0)=1$. Linear refers to the BIMP model without any nonlinearity. We find that using nonlinearities that meet the NOD criteria effectively prevent oversmoothing, while the others do not. We report the classification accuracy of our BIMP model with alternative nonlinearities in the NOD module in Table 10.
Table 10: Nonlinear opinion dynamics nonlinearity ablation. Classification accuracy of our BIMP model on the Cora dataset using various nonlinearities.
Layer tanh softsign arctan sigmoid ReLu GELU linear
1 69.96$\pm$1.45 63.25$\pm$1.73 64.05$\pm$1.58 60.36$\pm$1.39 64.17$\pm$1.72 62.45$\pm$1.57 77.52$\pm$1.44
2 75.00$\pm$1.50 68.34$\pm$2.03 72.62$\pm$2.40 63.20$\pm$1.63 67.21$\pm$2.18 67.06$\pm$2.25 81.92$\pm$0.85
4 79.93$\pm$1.41 72.37$\pm$1.50 76.91$\pm$1.82 65.54$\pm$1.80 73.16$\pm$1.44 72.40$\pm$1.60 82.08$\pm$1.34
8 82.21$\pm$1.26 77.05$\pm$1.67 79.84$\pm$1.05 63.91$\pm$2.13 77.76$\pm$1.47 77.56$\pm$1.49 81.24$\pm$1.48
16 82.83$\pm$1.12 79.88$\pm$1.82 81.81$\pm$1.54 29.55$\pm$1.80 81.32$\pm$0.96 81.47$\pm$0.63 80.45$\pm$1.43
32 82.81$\pm$1.19 81.45$\pm$1.48 82.48$\pm$1.81 29.92$\pm$1.22 82.51$\pm$1.51 27.99$\pm$4.22 79.99$\pm$1.21
64 82.53$\pm$1.07 81.14$\pm$1.65 82.94$\pm$0.73 30.72$\pm$1.02 76.56$\pm$3.88 29.64$\pm$1.82 76.74$\pm$1.86
128 82.18$\pm$1.06 81.71$\pm$1.37 81.26$\pm$1.93 29.37$\pm$2.77 71.18$\pm$7.09 26.27$\pm$4.89 75.44$\pm$0.89